# Simple-RF: Regularizing Sparse Input Radiance Fields with Simpler Solutions

### v4

@@ 1. Introduction
Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) show unprecedented levels of performance in synthesizing novel views of a scene by learning a volumetric representation implicitly within the weights of multi-layer perceptrons (MLP). Although NeRFs are very promising for view synthesis, there is a need to improve their design in a wide array of scenarios. For example, NeRFs have been enhanced to optimize and render quickly (Müller et al., 2022), reduce aliasing artifacts (Barron et al., 2021), and learn on unbounded scenes (Barron et al., 2022). Yet all these models require tens to hundreds of images per scene to learn the scene geometry accurately, and their quality deteriorates significantly when only a few training images are available (Jain et al., 2021). In this work, we focus on training both implicit radiance fields such as NeRF and explicit radiance fields such as TensoRF (Chen et al., 2022b) and ZipNeRF (Barron et al., 2023) with a sparse set of input images and aim to design novel regularizations for effective training.
@@
Researchers have extensively studied the training of NeRFs with sparse input views. One approach to training NeRFs with sparse input views is to use generalized NeRFs, where the NeRF is additionally conditioned on a latent scene representation obtained using a convolutional neural network. However, these models require a large multi-view dataset for pre-training and may suffer from generalization issues when used to render a novel scene (Niemeyer et al., 2022). The other thread of work on sparse input NeRFs follows the original NeRF paradigm of training scene-specific NeRFs, and designs novel regularizations to assist NeRFs in converging to a better scene geometry (Zhang et al., 2021b; Ni et al., 2024; Guo et al., 2024). One popular approach among such models is to supervise the depth estimated by the NeRF. RegNeRF (Niemeyer et al., 2022), DS-NeRF (Deng et al., 2022) and ViP-NeRF (Somraj and Soundararajan, 2023) use simple priors such as depth smoothness, sparse depth or relative depth respectively obtained through classical methods. On the other hand, DDP-NeRF (Roessle et al., 2022) and SCADE (Uy et al., 2023) pre-train convolutional neural networks (CNN) on a large dataset of scenes to learn a dense depth prior. These approaches may also suffer from issues similar to those of the generalized models. This raises the question of whether we can instead learn the dense depth supervision in-situ without employing any pre-training.
@@
Recently, there is also a growing interest in sparse input explicit radiance fields owing to their fast optimization and rendering times. However, the regularizations designed for these models are limited to a specific explicit radiance field and do not generalize to more recent models. For example, while the regularizations designed for ZeroRF (Shi et al., 2024) can be applied to TensoRF based models only, other regularizations are applicable to implicit models only (Yang et al., 2023; Zhu et al., 2024). It is desirable to design regularizations that are relevant to different radiance field models through a single framework. While there exists a plethora of implicit and explicit radiance field models, we consider the NeRF as the representative model for implicit radiance fields and consider two explicit radiance fields, namely, TensoRF (Chen et al., 2022b) and ZipNeRF (Barron et al., 2023). We note that although the NeRF based models are slow in optimization and rendering, NeRFs are versatile in learning different properties of the scenes  (Verbin et al., 2022; Bi et al., 2020; Srinivasan et al., 2021; Zhang et al., 2021a) and may be easier to optimize as compared to the explicit models. Hence, we believe that designing regularizations for sparse-input NeRF is also of considerable interest.
@@ 2.1. Implicit Radiance Fields
View Hallucination based Methods: Another line of regularization based approaches simulate dense sampling by hallucinating new viewpoints and regularizing the NeRF on different aspects such as semantic consistency (Jain et al., 2021), depth smoothness (Niemeyer et al., 2022), sparsity of mass (Kim et al., 2022) and depth based reprojection consistency (Kwak et al., 2023; Xu et al., 2022; Bortolon et al., 2022; Chen et al., 2022a). Instead of sampling new viewpoints randomly, FlipNeRF (Seo et al., 2023a) utilizes ray reflections to determine new viewpoints. Deceptive-NeRF (Liu et al., 2023) and ReconFusion (Wu et al., 2024) employ a diffusion model to generate images in hallucinated views and use the generated views in addition to the input views to train the NeRF. However, supervision with generative models could lead to content hallucinations, leading to poor fidelity (Lee and Lee, 2024).
Other regularizations: A few models also explore regularizations other than depth supervision and view hallucinations. FreeNeRF (Yang et al., 2023) and MI-MLP-NeRF (Zhu et al., 2024) regularize the NeRF by modifying the inputs. Specifically, FreeNeRF anneals the frequency range of positional encoded NeRF inputs as the training progresses, and MI-MLP-NeRF adds the 5D inputs to every layer of the NeRF MLP. MixNeRF (Seo et al., 2023b) models the volume density along a ray as a mixture of Laplacian distributions. Philip and Deschaintre (2023) scale the gradients corresponding to 3D points close to the camera when sampling the 3D points in inverse depth to reduce floaters close to the camera. VDN-NeRF (Zhu et al., 2023b) on the other hand, aims to resolve shape-radiance ambiguity in the case of dense input views. However, these approaches are designed for specific cases and are either sub-optimal or do not extend to more recent radiance field models.
@@ 2.2. Explicit Radiance Fields
The NeRF takes a long time to optimize and render novel views due to the need to query the NeRF MLP hundreds of times to render a single pixel. Hence, a common approach to fast optimization and rendering is to reduce the time taken per query. Early works such as PlenOctress (Yu et al., 2021a) and KiloNeRF (Reiser et al., 2021) focus on improving only the rendering time by baking the trained NeRF into an explicit structure such as Octrees or thousands of tiny MLPs. PlenOxels (Fridovich-Keil et al., 2022) and DVGO (Sun et al., 2022) reduce the optimization time by directly optimizing voxel grids, but at the cost of large memory requirements to store the voxel grids. TensoRF (Chen et al., 2022b) and K-Planes (Fridovich-Keil et al., 2023) reduce the memory consumption using factorized tensors that exploit the spatial correlation of the radiance field. Alternately, iNGP (Müller et al., 2022) and ZipNeRF (Barron et al., 2023) employ multi-resolution hash-grids to reduce the memory consumption. Recently, 3DGS (Kerbl et al., 2023) propose an alternative volumetric model for real-time rendering of novel views. Specifically, 3DGS employs 3D Gaussians to represent the scene and renders a view by splatting the Gaussians onto the corresponding image plane. While the above methods enable fast optimization and rendering, their performance still reduces significantly with fewer input views.
Recently, there is increasing interest in regularizing explicit models to learn with sparse inputs (Li et al., 2024; Yang et al., 2024). However, the regularizations designed in these models are limited to a specific explicit radiance field and do not generalize to other explicit models. For example, ZeroRF (Shi et al., 2024) imposes a deep image prior (Ulyanov et al., 2018) on the components of the TensoRF (Chen et al., 2022b) model. FSGS (Zhu et al., 2023a) and SparseGS (Xiong et al., 2023) improve the performance of 3DGS (Kerbl et al., 2023) in the sparse input case by improving the initialization of the 3D Gaussian point cloud and pruning Gaussians responsible for floaters respectively. On the other hand, our approach of regularizing with simpler solutions is applicable to various explicit models, such as TensoRF, iNGP and ZipNeRF as well as to implicit models such as NeRF.
@@ 4.1.4. Determining Reliable Depth Estimates
(8) $m_{a}=\begin{cases}1\qquad\ &\text{ if }(e_{a}\leq e_{m})\ \text{and}\ (e_{a}\leq e_{\tau})\\ 0&\text{ otherwise }.\end{cases}$
@@
[…]
@@ 4.1.5. Hierarchical Sampling
[…]
@@ 4.2. Simple-TensoRF
We first provide a brief overview of TensoRF (Chen et al., 2022b) in Sec. 4.2.1 and also describe the notation required to explain the design of our augmentations. We discuss the distortions observed with sparse input TensoRF in Sec. 4.2.2 and then discuss the design of augmentations in Sec. 4.2.3.
@@ 4.2.1. TensoRF Preliminaries
where $\gamma$ is the positional encoding described by Eq. 4. Thus, we note that $\mathcal{F}_{2}=\gamma\circ\mathcal{N}_{2}$. The color of the pixel is then obtained through volume rendering using $\sigma_{i}$ and $\mathbf{c}_{i}$ as in Sec. 3. For further details, we refer the readers to TensoRF (Chen et al., 2022b).
@@ 6.1.4. Ablations
We test the importance of each of the components of our model by disabling them one at a time. We disable the smoothing and Lambertian augmentations and coarse-fine consistency loss individually. When disabling ${\mathcal{L}_{cfc}}$, we additionally add augmentations to the fine NeRF since the knowledge learned by coarse NeRF may not efficiently propagate to the fine NeRF. We also analyze the need to supervise with only the reliable depth estimates by disabling the mask and stop-gradients in ${\mathcal{L}_{aug}}$ and ${\mathcal{L}_{cfc}}$. In addition, we also analyze the effect of including residual positional encodings $\gamma(\mathbf{p}_{i},l_{p}^{s},l_{p})$ while predicting the color in the smoothing augmentation model. Tab. 5 shows a quantitative comparison between the ablated models. We observe that each of the components is crucial, and disabling any of them leads to a drop in performance. Further, using all the depths for supervision instead of only the reliable depths leads to a significant drop in performance. Finally, disabling ${\mathcal{L}_{cfc}}$ also leads to a drop in performance in addition to increasing the training time by almost 2×2$\times$2 × due to the inclusion of augmentations for the fine NeRF.
@@ 6.2.1. Implementation Details
Building on the original TensoRF code base, we employ Adam Optimizer with an initial learning rate of $22e-$ 2 and $31e-$ 3 for the tensor and MLP parameters respectively, which decay to $32e-$ 3 and $41e-$ 4. We employ the same hyper-parameters as the original implementation for the main model as follows: $R_{\sigma}=24$, $R_{c}=72$, $\mathbf{b}=\left\{(-1.5,1.5),(-1.67,1.67),(-1.0,1.0)\right\}$, ${N_{vox}}=640^{3}$, $D=27$, and $l_{v}=0$. We set $R_{\sigma}^{s}=12$, $b_{z_{1}}^{s}=-0.5$ and ${N_{vox}^{s}}=160^{3}$, ${N_{mc}}=5$, $k=5$, $e_{\tau}=0.1$ for the augmented model and the remaining hyper-parameters are the same as the main model. We weigh the losses as ${\lambda_{m}}={\lambda_{a}}=1,{\lambda_{sd}}={\lambda_{aug}}=0.1,{\lambda_{mc}}=0.01\text{ and }{\lambda_{cfc}}=0$. We train the models on a single NVIDIA RTX 2080 Ti 11GB GPU for 25k iterations and enable ${\mathcal{L}_{aug}}$ after 5k iterations.
@@ 6.3. Simple-ZipNeRF
Table 7. Quantitative results of ZipNeRF based models on the MipNeRF360 dataset.
@@ 6.3.2. Quantitative and Qualitative Results
We compare the performance of ZipNeRF with and without our augmentations on the MipNeRF360 and NeRF-Synthetic datasets in Tabs. 7 and 8 respectively. We observe that including our augmentations improves performance significantly on both datasets in terms of all the evaluation measures. This observation is further supported by the qualitative examples in Figs. 21, 22, 24 and 23, where we observe a clear improvement in the quality of the rendered novel views and depth when employing our augmentations. From Tab. 7, we observe that the performace of InstantSplat decreases with more number of input views. This is perhaps because of applying DUSt3R (Wang et al., 2024b) independently on each pair of input views, leading to accumulation of errors. While CoR-GS achieves very good performance, we find that it does not reconstruct high frequency details in the scene as seen in Fig. 22. As a result, it achieves better PSNR and SSIM scores, but poorer LPIPS scores. Further, we observe that CoR-GS also suffers from large floater artifacts as shown in Fig. 21(a). In addition, Tabs. 7 and 25 also shows the performance of the augmented model on the MipNeRF360 dataset. We observe a significant reduction in distortions in the renders of the augmented model; however, the same does not reflect in the quantitative evaluation due to the blur introduced by the augmented model.

### v5

@@ 1. Introduction
Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) show unprecedented levels of performance in synthesizing novel views of a scene by learning a volumetric representation implicitly within the weights of multi-layer perceptrons (MLP). Although NeRFs are very promising for view synthesis, there is a need to improve their design in a wide array of scenarios. For example, NeRFs have been enhanced to optimize and render quickly (Müller et al., 2022), reduce aliasing artifacts (Barron et al., 2021), and learn on unbounded scenes (Barron et al., 2022). Yet all these models require tens to hundreds of images per scene to learn the scene geometry accurately, and their quality deteriorates significantly when only a few training images are available (Jain et al., 2021). In this work, we focus on training both implicit radiance fields such as NeRF and explicit radiance fields such as TensoRF (Chen et al., 2022a) and ZipNeRF (Barron et al., 2023) with a sparse set of input images and aim to design novel regularizations for effective training.
@@
Researchers have extensively studied the training of NeRFs with sparse input views. One approach to training NeRFs with sparse input views is to use generalized NeRFs, where the NeRF is additionally conditioned on a latent scene representation obtained using a convolutional neural network. However, these models require a large multi-view dataset for pre-training and may suffer from generalization issues when used to render a novel scene (Niemeyer et al., 2022). The other thread of work on sparse input NeRFs follows the original NeRF paradigm of training scene-specific NeRFs, and designs novel regularizations to assist NeRFs in converging to a better scene geometry (Zhang et al., 2021a; Ni et al., 2024; Guo et al., 2024). One popular approach among such models is to supervise the depth estimated by the NeRF. RegNeRF (Niemeyer et al., 2022), DS-NeRF (Deng et al., 2022) and ViP-NeRF (Somraj and Soundararajan, 2023) use simple priors such as depth smoothness, sparse depth or relative depth respectively obtained through classical methods. On the other hand, DDP-NeRF (Roessle et al., 2022) and SCADE (Uy et al., 2023) pre-train convolutional neural networks (CNN) on a large dataset of scenes to learn a dense depth prior. These approaches may also suffer from issues similar to those of the generalized models. This raises the question of whether we can instead learn the dense depth supervision in-situ without employing any pre-training.
@@
Recently, there is also a growing interest in sparse input explicit radiance fields owing to their fast optimization and rendering times. However, the regularizations designed for these models are limited to a specific explicit radiance field and do not generalize to more recent models. For example, while the regularizations designed for ZeroRF (Shi et al., 2024) can be applied to TensoRF based models only, other regularizations are applicable to implicit models only (Yang et al., 2023; Zhu et al., 2024). It is desirable to design regularizations that are relevant to different radiance field models through a single framework. While there exists a plethora of implicit and explicit radiance field models, we consider the NeRF as the representative model for implicit radiance fields and consider two explicit radiance fields, namely, TensoRF (Chen et al., 2022a) and ZipNeRF (Barron et al., 2023). We note that although the NeRF based models are slow in optimization and rendering, NeRFs are versatile in learning different properties of the scenes  (Verbin et al., 2022; Bi et al., 2020; Srinivasan et al., 2021; Zhang et al., 2021b) and may be easier to optimize as compared to the explicit models. Hence, we believe that designing regularizations for sparse-input NeRF is also of considerable interest.
@@ 2.1. Implicit Radiance Fields
View Hallucination based Methods: Another line of regularization based approaches simulate dense sampling by hallucinating new viewpoints and regularizing the NeRF on different aspects such as semantic consistency (Jain et al., 2021), depth smoothness (Niemeyer et al., 2022), sparsity of mass (Kim et al., 2022) and depth based reprojection consistency (Kwak et al., 2023; Xu et al., 2022; Bortolon et al., 2022; Chen et al., 2022b). Instead of sampling new viewpoints randomly, FlipNeRF (Seo et al., 2023a) utilizes ray reflections to determine new viewpoints. Deceptive-NeRF (Liu et al., 2023) and ReconFusion (Wu et al., 2024) employ a diffusion model to generate images in hallucinated views and use the generated views in addition to the input views to train the NeRF. However, supervision with generative models could lead to content hallucinations, leading to poor fidelity (Lee and Lee, 2024).
Other regularizations: A few models also explore regularizations other than depth supervision and view hallucinations. FreeNeRF (Yang et al., 2023) and MI-MLP-NeRF (Zhu et al., 2024) regularize the NeRF by modifying the inputs. Specifically, FreeNeRF anneals the frequency range of positional encoded NeRF inputs as the training progresses, and MI-MLP-NeRF adds the 5D inputs to every layer of the NeRF MLP. MixNeRF (Seo et al., 2023b) models the volume density along a ray as a mixture of Laplacian distributions. Philip and Deschaintre (2023) scale the gradients corresponding to 3D points close to the camera when sampling the 3D points in inverse depth to reduce floaters close to the camera. VDN-NeRF (Zhu et al., 2023a) on the other hand, aims to resolve shape-radiance ambiguity in the case of dense input views. However, these approaches are designed for specific cases and are either sub-optimal or do not extend to more recent radiance field models.
@@ 2.2. Explicit Radiance Fields
The NeRF takes a long time to optimize and render novel views due to the need to query the NeRF MLP hundreds of times to render a single pixel. Hence, a common approach to fast optimization and rendering is to reduce the time taken per query. Early works such as PlenOctress (Yu et al., 2021a) and KiloNeRF (Reiser et al., 2021) focus on improving only the rendering time by baking the trained NeRF into an explicit structure such as Octrees or thousands of tiny MLPs. PlenOxels (Fridovich-Keil et al., 2022) and DVGO (Sun et al., 2022) reduce the optimization time by directly optimizing voxel grids, but at the cost of large memory requirements to store the voxel grids. TensoRF (Chen et al., 2022a) and K-Planes (Fridovich-Keil et al., 2023) reduce the memory consumption using factorized tensors that exploit the spatial correlation of the radiance field. Alternately, iNGP (Müller et al., 2022) and ZipNeRF (Barron et al., 2023) employ multi-resolution hash-grids to reduce the memory consumption. Recently, 3DGS (Kerbl et al., 2023) propose an alternative volumetric model for real-time rendering of novel views. Specifically, 3DGS employs 3D Gaussians to represent the scene and renders a view by splatting the Gaussians onto the corresponding image plane. While the above methods enable fast optimization and rendering, their performance still reduces significantly with fewer input views.
Recently, there is increasing interest in regularizing explicit models to learn with sparse inputs (Li et al., 2024; Yang et al., 2024). However, the regularizations designed in these models are limited to a specific explicit radiance field and do not generalize to other explicit models. For example, ZeroRF (Shi et al., 2024) imposes a deep image prior (Ulyanov et al., 2018) on the components of the TensoRF (Chen et al., 2022a) model. FSGS (Zhu et al., 2023b) and SparseGS (Xiong et al., 2023) improve the performance of 3DGS (Kerbl et al., 2023) in the sparse input case by improving the initialization of the 3D Gaussian point cloud and pruning Gaussians responsible for floaters respectively. On the other hand, our approach of regularizing with simpler solutions is applicable to various explicit models, such as TensoRF, iNGP and ZipNeRF as well as to implicit models such as NeRF.
@@ 4.1.4. Determining Reliable Depth Estimates
[…]
@@
(9) ${\mathcal{L}_{aug}}=\mathbbm{1}_{\{m_{a}=1\}}\odot\|z_{m}-\cancel{\nabla}(z_{a})\|^{2}+\mathbbm{1}_{\{m_{m}=1\}}\odot\|\cancel{\nabla}(z_{m})-z_{a}\|^{2},$
@@ 4.1.5. Hierarchical Sampling
(10) ${\mathcal{L}_{cfc}}=\mathbbm{1}_{\{m_{f}=1\}}\odot\|z_{c}-\cancel{\nabla}(z_{f})\|^{2}+\mathbbm{1}_{\{m_{c}=1\}}\odot\|\cancel{\nabla}(z_{c})-z_{f}\|^{2},$
@@ 4.2. Simple-TensoRF
We first provide a brief overview of TensoRF (Chen et al., 2022a) in Sec. 4.2.1 and also describe the notation required to explain the design of our augmentations. We discuss the distortions observed with sparse input TensoRF in Sec. 4.2.2 and then discuss the design of augmentations in Sec. 4.2.3.
@@ 4.2.1. TensoRF Preliminaries
where $\gamma$ is the positional encoding described by Eq. 4. Thus, we note that $\mathcal{F}_{2}=\gamma\circ\mathcal{N}_{2}$. The color of the pixel is then obtained through volume rendering using $\sigma_{i}$ and $\mathbf{c}_{i}$ as in Sec. 3. For further details, we refer the readers to TensoRF (Chen et al., 2022a).
@@ 6.1.4. Ablations
We test the importance of each of the components of our model by disabling them one at a time. We disable the smoothing and Lambertian augmentations and coarse-fine consistency loss individually. When disabling $\mathcal{L}_{cfc}$}, we additionally add augmentations to the fine NeRF since the knowledge learned by coarse NeRF may not efficiently propagate to the fine NeRF. We also analyze the need to supervise with only the reliable depth estimates by disabling the mask and stop-gradients in $\mathcal{L}_{aug}$} and $\mathcal{L}_{cfc}$}. In addition, we also analyze the effect of including residual positional encodings $\gamma(\mathbf{p}_{i},l_{p}^{s},l_{p})$ while predicting the color in the smoothing augmentation model. Tab. 5 shows a quantitative comparison between the ablated models. We observe that each of the components is crucial, and disabling any of them leads to a drop in performance. Further, using all the depths for supervision instead of only the reliable depths leads to a significant drop in performance. Finally, disabling $\mathcal{L}_{cfc}$} also leads to a drop in performance in addition to increasing the training time by almost $2\times$ due to the inclusion of augmentations for the fine NeRF.
@@ 6.2.1. Implementation Details
Building on the original TensoRF code base, we employ Adam Optimizer with an initial learning rate of 2e22e-2 and 1e31e-3 for the tensor and MLP parameters respectively, which decay to 2e32e-3 and 1e41e-4. We employ the same hyper-parameters as the original implementation for the main model as follows: $R_{\sigma}=24$, $R_{c}=72$, $\mathbf{b}=\left\{(-1.5,1.5),(-1.67,1.67),(-1.0,1.0)\right\}$, ${N_{vox}}=640^{3}$, $D=27$, and $l_{v}=0$. We set $R_{\sigma}^{s}=12$, $b_{z_{1}}^{s}=-0.5$ and ${N_{vox}^{s}}=160^{3}$, ${N_{mc}}=5$, $k=5$, $e_{\tau}=0.1$ for the augmented model and the remaining hyper-parameters are the same as the main model. We weigh the losses as ${\lambda_{m}}={\lambda_{a}}=1,{\lambda_{sd}}={\lambda_{aug}}=0.1,{\lambda_{mc}}=0.01\text{ and }{\lambda_{cfc}}=0$. We train the models on a single NVIDIA RTX 2080 Ti 11GB GPU for 25k iterations and enable $\mathcal{L}_{aug}$} after 5k iterations.
@@ 6.3. Simple-ZipNeRF
Table 7. Quantitative results of ZipNeRF and 3DGS based models on the MipNeRF360 dataset.
@@ 6.3.2. Quantitative and Qualitative Results
We compare the performance of ZipNeRF with and without our augmentations on the MipNeRF360 and NeRF-Synthetic datasets in Tabs. 7 and 8 respectively. We observe that including our augmentations improves performance significantly on both datasets in terms of all the evaluation measures. This observation is further supported by the qualitative examples in Figs. 21, 22, 24 and 23, where we observe a clear improvement in the quality of the rendered novel views and depth when employing our augmentations. From Tab. 7, we observe that the performace of InstantSplat decreases with more number of input views. This is perhaps because of applying DUSt3R (Wang et al., 2024a) independently on each pair of input views, leading to accumulation of errors. While CoR-GS achieves very good performance, we find that it does not reconstruct high frequency details in the scene as seen in Fig. 22. As a result, it achieves better PSNR and SSIM scores, but poorer LPIPS scores. Further, we observe that CoR-GS also suffers from large floater artifacts as shown in Fig. 21(a). In addition, Tabs. 7 and 25 also shows the performance of the augmented model on the MipNeRF360 dataset. We observe a significant reduction in distortions in the renders of the augmented model; however, the same does not reflect in the quantitative evaluation due to the blur introduced by the augmented model.
@@ 7.1. Computational Complexity
7.2. Analysis
(a) NeRF-LLFF Dataset
(b) RealEstate10K Dataset
Figure 28. Correlation between spatial complexity and the SSIM scores of Simple-NeRF across multiple scenes and number of input views. We observe a strong negative correlation i.e., the performance of Simple-NeRF is very good for scenes with low spatial complexity and the performance reduces with increasing spatial complexity.
Figure 29. Performance improvement of Simple-NeRF over DS-NeRF baseline with respect to the baseline performance. We observe that Simple-NeRF provides significant improvement when the baseline performance of ZipNeRF is moderate. When the baseline performance is either very poor or very good, the improvement is marginal.
We now analyse the correlation between the performance of our model against the scene complexity and the performance of the base model. We measure the spatial complexity of a scene using the average image gradient magnitude across all the training images, normalized by the number of input views. The intuition is that our model is good at reducing the floaters in smooth regions, but struggles in highly textured regions as they are more difficult to reconstruct accurately with sparse views. However, as the number of input views increases, our model is able to reconstruct more complex scenes. In Figure 28, we show a scatter plot between the spatial complexity and the SSIM scores of Simple-NeRF across multiple scenes in NeRF-LLFF and RealEstate10K datasets with 2, 3 and 4 input views. We observe that Simple-NeRF achieves a SSIM in excess of 0.8 for scenes with spatial complexity less than $6$ across both the datasets. Further, the plots provide a coarse trade-off between the spatial complexity and the achievable SSIM score for a given number of input views. Thus, one could determine whether our model is likely to achieve a target SSIM score for a given scene by computing its spatial complexity.
Next, we analyse the dependence of the improvement achieved by Simple-NeRF over the base NeRF model given its results. Specifically, we find that Simple-NeRF achieves a higher improvement when the base model has a moderate performance, but still suffers from floaters. We quantify this observation through the following analysis. We hypothesize that the improvement achieved by Simple-NeRF depends on both the SSIM score and the amount of floaters in the base model. Mathematically, if $s_{b}$ is the SSIM of the base model and $f_{b}$ is the fraction of pixels corresponding to floaters in the base model, we show a scatter plot between $\alpha s_{b}-\beta f_{b}$ and the improvement in SSIM in Figure 29, where $\alpha$ and $\beta$ are normalization constants. We subtract $f_{b}$ from $s_{b}$, since they are complementary in nature. We determine the floater pixels as those whose depth is less than half of the median depth of the scene. We set $\alpha=0.2$ and $\beta=0.5$ through a simple grid search. We observe that the improvement achieved by Simple-NeRF increases as the base model performance improves and the amount of floaters decreases. The quantum of improvement reduces as the base model performance saturates and the amount of floaters reduces.